Simulation production news

Speaker : Jason Webb ( BNL )


Talk time : 12:20, Duration : 00:15

Update on LFSpectra Request

30M events were requested, 5000 jobs satisfy this request.  On Friday we ramped down production (with sufficient number of jobs to complete the request) in preparation for the cataclysm (Irene).  Tuesday Lidia and Leve discovered that 700 jobs had duplicate IDs, thus 700 jobs were overwritten when copied to local directories for staging to HPSS and disk.  Yesterday, 700 jobs were started locally.  680/700 jobs completed this morning, yielding a total of 4980/5000 jobs completed in this request.  We are declaring this a success, and pushing files to HPSS, disk and cataloging.

o Note: for future requests which are run at multiple sites and/or by multple deputies... jobs should be submitted with large offsets in job ID to avoid such clashes.

Update on Sti/Stv/CA Tracker Evalulation Samples

Production at PDSF

Staged to /star/data16/simu/rcf11020 .. 11025

11020 (simple events, 25 muons/event) done.
11021 W+/- w/ pileup beginning
others to follow.  Est 1 week wall time for initial samples.
 

Next up: pp 62 GeV minbias sample 200k